OpenStudy (anonymous):

translate into an equation and solve. the perimeter of a rectangular driveway is 290 ft. the length is 1ft longer than twice the width. find the dimensions of the driveway. show work please

OpenStudy (kropot72):

Let the width of the driveway = w feet Then the length = (2w + 1) The perimeter of a rectangle = (2 * width) + (2 * length) Can you find the equation for the perimeter of the driveway in question now?

OpenStudy (anonymous):

not exactly, hang on

OpenStudy (kropot72):

Perimeter = 290 = 2w + 2(2w + 1)

OpenStudy (anonymous):

so just find w?

OpenStudy (kropot72):

Yes, just solve for w.

OpenStudy (anonymous):

ooooh ok !!! now i can get it then

OpenStudy (anonymous):

290=2w+2(2w+1) 290=2w+4w+2 290=6w+2 290= what do i do next?

OpenStudy (anonymous):

i know im trying to get w on its own but wheich side do i start on?

OpenStudy (kropot72):

Good work! Now just subtract 2 from each side of the equation to get an expression for 6w.

OpenStudy (anonymous):

288=6w then just divide each side by 6?to get w alone?

OpenStudy (kropot72):

Correct. Then just double the result and add 1 to find the length in feet.

OpenStudy (anonymous):

w=48

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

97? because to double i do 48*2=96+1=97

OpenStudy (kropot72):

Well done :) Your results for width and length are correct. Don't forget to add the units (feet) to your answer.

OpenStudy (anonymous):

GOT IT THANKS!!

OpenStudy (kropot72):

You're welcome :)